# Stochastic Re-weighted Gradient Descent via Distributionally Robust Optimization

### v4

@@ B.3 Other Divergences
Appendix C DomainBed Benchmark
Appendix D Choice of divergence in RGD
@@
Appendix E Comparisons against KL DRO optimization approaches
In this section, we discuss a few more aspects of the related work that weren’t discussed in the main paper. We specifically focus on prior works that developed algorithms for KL divergence based DRO. The earliest work on KL-DRO dates back to 2013 by Hu & Hong . However, it was only recently that these works have become widespread in deep learning. The RECOVER algorithm by Qi et al. (2021) was one of the early works to scale KL-DRO to deep neural networks. It attempted to solve the non-convex DRO problem with a duality-free stochastic method by formulating the min-max formulation into an equivalent stochastic compositional problem. ABSGD (Qi et al., 2020) and SCDRO (Qi et al., 2022) improved upon RECOVER by designing more efficient algorithms. However, the performance of these algorithms on large-scale models and datasets is not rigorously studied, as the Imagenet-LT, iNaturalist experiments conducted in these works started from a pre-trained network and only finetuned the last layer. In contrast, in this work, we learn the entire ViT-S model from scratch and show improved generalization.
Appendix F Additional Experimental Results and Missing Details
This section provides additional experimental results and details that are missing in the main paper. The search space and hyperparameters tuned in our experiments of are depicted in Appendix F.1
F.1 Hyperparameter Tuning
@@
F.2 Toy Examples
In this section, we perform two simple experiments to better understand the robustness properties of our proposed approach.
@@ F.2 Toy Examples
Class Imbalance and Fairness: Figure 5 compares RGD with class-imbalanced loss and focal loss. Observe that our method outperforms these baselines by a significant margin of $2$-$4\%$ (see Figure 5). Table 4 compares RGD with ABSGD, TERM. The ABSGD numbers in the table are from Qi et al. (2020) and are obtained by running the optimization procedure for significantly larger number of epochs (200) in comparison to RGD which was only run for 120 epochs. Observe that RGD outperforms both ABSGD and TERM by $\approx 1\%$, on average. Refer to Appendix F for more details about the experiment.
@@
F.3 Ablations and Comparison with prior KL-DRO techniques
For TERM, we use the batch (non-hierarchical) version (as shown in Algorithm 1 of (Li et al., 2023)) - requiring two degrees of hyperparameters (tilting coefficient $t$ and the learning rate ($lr$)). For the tilting coefficient, we use a search space of (Li et al., 2020): $\{0.2,0.5,1,3,5\}$. For the learning rate, we use a lr multiplier to the baseline run as $\{0.7,0.8,0.9,1,1.1,1.2,1.3\}$. For the stochastic version of TERM, which is identical to ABSGD (Qi et al., 2020) baseline - requires an additional coefficient of moving average ($\beta$). We use a grid search space of $\{0.25,0.5,0.75\}$ for tuning $\beta$, and tune the learning rate in a similar fashion to TERM. We also tune $\lambda$ (similar to the tilting coefficient of TERM) in the search space $\{1,3,5,7\}$.
F.4 Vanilla Classification
This section briefly discusses a few additional results from our experiments on Vanilla classification. Table 10 depicts the performance of our other variant RGD-RevKL in comparison to the baseline approach. Furthermore, we also show fine-tuning improvements of EfficientNet-v2-l over various tasks such as Cars and Food101 as depicted in Table 11. Furthermore, we also demonstrate that our proposed approach is simple and shows significant improvements, not only for SOTA approaches but also basic MLP procedures as depicted in Table 16, Table 17, and Table 18. These tables help showcase that the simple addition of our proposed approach does show significant improvements of +2.77% (in accuracy) on multi-class and +1.77% (in AUROC) on binary class tasks respectively.
bert-base-uncased MNLI QQP QNLI SST-2 MRPC RTE COLA Avg
Default 81.33 89.62 87.93 90.63 89.55 67.19 54.53 80.11
RGD-RevKL (Ours) 82.97 89.87 90.79 91.28 88.54 71.23 59.28 81.97
RGD (Ours) 83.06 91.06 90.35 91.78 88.28 71.48 58.56 82.05
Table 11: Ablation study to showcase the gain achieved by using RGD for various tasks in Vision benchmarks for ViT-S on Imagenet-1K and EfficientNet-v2-l on Finetuning tasks.
Imagenet-1K Cars-FineTuning Food101-FineTuning
Default 78.1 92.03 92.65
RGD (Ours) 79.0 92.62 92.75
Furthermore, we extend our work to large-scale tasks in NLP such as LLM pre-training which has become more prevalent over the recent years. Preliminary results on miniGPT pre-training as illustrated in this section showcase the efficacy of our approach in these settings.
miniGPT (Zhu et al., 2023) - is a minimal implementation of a decoder-only transformer language model. We consider a 6 layer model and train on the lm1B small dataset which has 1B tokens. We trained for 100K steps with a batch size of 256. We used the default learning rate of 0.0016 for the baseline. For RGD, we fix the clipping threshold to 1 and tune the learning rate. We achieved 1% improvement on the eval perplexity score. Table 12 illustrates our results in this setting.
Table 12: Results on standard Lm1b pre-training task with miniGPT. Off-the-hat addition of RGD  leads to improvements of evaluation perplexity by $1\%$.
Dataset Lm1b eval Perplexity
Default 2.9218
RGD (Ours) 2.8979
F.5 Class Imbalance Experiments
@@
This section briefly discusses additional results from our experiments on the Class Imbalance domain with datasets such as CIFAR-10 and CIFAR-100. It is well known that DRO outputs models with good tail performance (Duchi & Namkoong, 2018). Since RGD directly solves the DRO objective, our models are also naturally endowed with this property. To demonstrate this, we extend our experiments on linear regression to a more realistic image dataset, where some classes appear very rarely in the data set while some appear very frequently. We use the Long-Tailed CIFAR dataset, where we reduce the number of training samples per class according to an exponential function as proposed by (Cui et al., 2019). We define the imbalance factor of a dataset as the number of training samples in the largest class divided by the smallest. Similar to the works of (Shu et al., 2019), we use a ResNet-32 architecture for training. Apart from Cross Entropy loss, we also include Focal Loss (Lin et al., 2017) and Class Balanced Loss (Cui et al., 2019) as additional baselines. We also experimented with the long-tailed CIFAR-100 dataset and showed that our proposed approach could again show significant improvements. Figure 5 illustrates the performance of our approach in comparison to other state-of-the-art methods. Overall, in comparison to the SOTA approach in this task (Class Balanced Loss), our proposed approach brings about an improvement of +0.79%. A more comprehensive comparison with additional state-of-the-art baselines such as L2RW (Ren et al., 2018), and Meta-Weight-Net (Shu et al., 2019) is illustrated in Table 14. Although these models use additional data as a meta-validation-set, our proposed approach outperforms L2RW and is roughly competitive with the Meta-Weight-Net model.
@@
Table 13 depicts the accuracy metric of models on various levels of the imbalance factor. From Table 13, we show that our proposed approach RGD outperforms other baselines such as Focal Loss and Class Balanced Loss by +0.79%. Furthermore, when models are trained on additional data, either by fine-tuning or by using a meta-learning framework to learn weights (such as Meta-Weight-Net and L2RW), we show that our proposed approach is competitively similar (-0.22%). Table 14 illustrates this analysis further. The performance metrics of the baseline approaches were taken from (Shu et al., 2019). Additional comparisons against other losses, such as cross-entropy with label smoothing and large margin softmax loss, are shown in Table 15.
@@
Focal Loss (Lin et al., 2017) 65.29 70.38 76.71 82.76 86.66 93.03 79.14 35.62 38.41 44.32 51.95 55.78 70.52 49.43
Class Balanced Loss (Cui et al., 2019) 68.89 74.57 79.27 84.36 87.49 92.89 81.25 36.23 39.60 45.32 52.59 57.99 70.50 50.21
@@
L2RW (Ren et al., 2018) $\star$ 66.51 74.16 78.93 82.12 85.19 89.25 77.69 33.38 40.23 44.44 51.64 53.73 64.11 47.92
Meta-Weight-Net (Shu et al., 2019) $\star$ 68.91 75.21 80.06 84.94 87.84 92.66 81.60 37.91 42.09 46.74 54.37 58.46 70.37 51.65
@@
F.6 Tabular Classification
This section discusses a few additional results from our experiments on Tabular classification. Table 19 depicts our proposed approach’s accuracy compared to other baselines on multi-class tabular datasets. Our method outperforms previous SOTA in this problem by +1.27%. Furthermore, Table 20 illustrates the AUROC score of our proposed approach in comparison to state-of-the-art baselines on binary-class tabular datasets. Our approach shows an improvement of +1.5% in this setting as well. The performance metrics of the baseline approaches were taken from (Majmundar et al., 2022).
@@ F.6 Tabular Classification
RF (Breiman, 2001) 88.43 42.73 97.62 71.37 75.04
GBDT (Friedman, 2001) 88.71 45.7 100 72.96 76.84
RF-G (Rahimi & Recht, 2008) 89.84 29.32 97.65 71.57 72.10
MET-R (Majmundar et al., 2022) 88.84 28.94 97.44 69.68 71.23
VIME (Yoon et al., 2020) 80.36 34.00 95.77 62.80 68.23
DACL+ (Verma et al., 2021) 81.40 39.70 91.40 64.23 69.18
SubTab (Ucar et al., 2021) 87.59 39.34 98.31 42.36 66.90
TabNet (Arik & Pfister, 2019) 88.18 33.75 96.63 65.13 70.92
MET (Majmundar et al., 2022) 91.68 47.82 99.19 76.71 78.85
@@
RF (Breiman, 2001) 64.36 91.53 77.57 99.62 83.27
GBDT (Friedman, 2001) 64.4 92.5 78.77 99.34 83.75
RF-G (Rahimi & Recht, 2008) 54.45 90.09 80.32 52.65 69.37
MET-R (Majmundar et al., 2022) 53.2 83.54 79.17 82.03 74.49
VIME (Yoon et al., 2020) 57.27 87.37 74.28 94.87 78.45
DACL+ (Verma et al., 2021) 61.18 89.01 75.32 86.63 78.04
SubTab (Ucar et al., 2021) 64.92 88.95 76.57 88.93 79.00
TabNet (Arik & Pfister, 2019) 69.40 77.30 80.91 96.98 81.15
@@
F.7 Meta-Learning
This section discusses some additional results from our experiments in the meta-learning domain. Table 21 depicts a complete table and comparison of our proposed approach on the MAML baseline compared to others. Overall, we notice improvement across the board, especially in the outliers, as shown in the Worst-kmetrics. Note that although our RGD has been applied on MAML (Finn et al., 2017) in our current experiments, our approach is analogous to the model and can be extended to other meta-learning techniques such as Protonet (Snell et al., 2017), CNAPs (Requeima et al., 2019), etc. as well. The performance metrics of the baseline approaches were taken from (Kumar et al., 2022).
@@ F.8 DomainBed
ERM (Gulrajani & Lopez-Paz, 2020) 85.5 $\pm$ 0.1 77.5 $\pm$ 0.4 66.5 $\pm$ 0.2 40.9 $\pm$ 0.1 67.6
IRM (Arjovsky et al., 2019) 83.5 $\pm$ 0.8 78.5 $\pm$ 0.5 64.3 $\pm$ 2.2 33.9 $\pm$ 2.8 65.1
GroupDRO (Sagawa et al., 2019) 84.4 $\pm$ 0.8 76.7 $\pm$ 0.6 66.0 $\pm$ 0.7 33.3 $\pm$ 0.2 65.1
Mixup (Yan et al., 2020a) 84.6 $\pm$ 0.6 77.4 $\pm$ 0.6 68.1 $\pm$ 0.3 39.2 $\pm$ 0.1 67.33
MLDG (Li et al., 2018a) 84.9 $\pm$ 1.0 77.2 $\pm$ 0.4 66.8 $\pm$ 0.6 41.2 $\pm$ 0.1 67.53
CORAL (Sun & Saenko, 2016) 86.2$\pm$ 0.3 78.8$\pm$ 0.6 68.7$\pm$ 0.3 41.5 $\pm$ 0.1 68.8
MMD (Li et al., 2018b) 84.6 $\pm$ 0.5 77.5 $\pm$ 0.9 66.3 $\pm$ 0.1 23.4 $\pm$ 9.5 62.95
DANN (Ganin et al., 2016) 83.6 $\pm$ 0.4 78.6 $\pm$ 0.4 65.9 $\pm$ 0.6 38.3 $\pm$ 0.1 66.6
CDANN (Li et al., 2018c) 82.6 $\pm$ 0.9 77.5 $\pm$ 0.1 65.8 $\pm$ 1.3 38.3 $\pm$ 0.3 66.05
MTL (Blanchard et al., 2021) 84.6 $\pm$ 0.5 77.2 $\pm$ 0.4 66.4 $\pm$ 0.5 40.6 $\pm$ 0.1 67.2
SagNet (Nam et al., 2021) 86.3 $\pm$ 0.2 77.8 $\pm$ 0.5 68.1 $\pm$ 0.1 40.3 $\pm$ 0.1 68.13
ARM (Zhang et al., 2021) 85.1 $\pm$ 0.4 77.6 $\pm$ 0.3 64.8 $\pm$ 0.3 35.5 $\pm$ 0.2 65.75
VREx (Krueger et al., 2021) 84.9 $\pm$ 0.6 78.3 $\pm$ 0.2 66.4 $\pm$ 0.6 33.6 $\pm$ 2.9 65.8
RSC (Huang et al., 2020) 85.2 $\pm$ 0.9 77.1 $\pm$ 0.5 65.5 $\pm$ 0.9 38.9 $\pm$ 0.5 66.68
MIRO (Cha et al., 2022) 85.4 $\pm$ 0.4 79.0 $\pm$ 0.0 70.5 $\pm$ 0.4 44.3 $\pm$ 0.2 69.8
@@
In this section, we briefly discuss additional results from our DomainBed experiments. Table 22 depicts a complete table and comparison of our proposed approach to a multitude of state-of-the-art approaches in this field. Furthermore, we also show that our proposed approach outperforms previous SOTA by +0.7%. Furthermore, we also present the per-environment breakdown of our approach in various datasets in Table 23, Table 24, Table 25, and Table 26 for PACS, VLCS, OfficeHome, and DomainNet respectively. The performance metrics of the baseline approaches were taken from (Gulrajani & Lopez-Paz, 2020).
F.9 Convergence of RGD and additional costs

### v5

@@ B.3 Other Divergences
Appendix C Choice of divergence in RGD
@@
Appendix D Additional Experimental Results and Missing Details
This section provides additional experimental results and details that are missing in the main paper. The search space and hyperparameters tuned in our experiments of are depicted in Appendix D.1
D.1 Hyperparameter Tuning
@@
D.1.1 Existing KL-DRO techniques
For TERM, we use the batch (non-hierarchical) version (as shown in Algorithm 1 of Li et al. (2023)) - requiring two degrees of hyperparameters (tilting coefficient $t$ and the learning rate ($lr$)). For the tilting coefficient, we use a search space of Li et al. (2021): $\{0.2,0.5,1,3,5\}$. For the learning rate, we use a lr multiplier to the baseline run as $\{0.7,0.8,0.9,1,1.1,1.2,1.3\}$. For the stochastic version of TERM, which is identical to ABSGD (Qi et al., 2023b) baseline - requires an additional coefficient of moving average ($\beta$). We use a grid search space of $\{0.25,0.5,0.75\}$ for tuning $\beta$, and tune the learning rate in a similar fashion to TERM. We also tune $\lambda$ (similar to the tilting coefficient of TERM) in the search space $\{1,3,5,7\}$.
D.2 Toy Experiments
In this section, we perform a simple experiment to better understand the robustness properties of our proposed approach.
@@ D.2 Toy Experiments
D.3 Comparisons against Sharpness Aware Minimization
Below we compare with Sharpness-Aware Minimization (SAM) (Foret et al., 2021) both from theoretical and empirical perspectives. From a theoretical perspective, SAM performs robust optimization in the weight space (that is SAM tries to learn a model that is robust to perturbations of weights). In contrast, RGD performs robust optimization in the distribution space. So, RGD and SAM are orthogonal to each other and can potentially be merged together to boost the performance. In Table 11, we compare performance of various approaches for CIFAR-10 optimization. It can be seen that RGD marginally boosts the performance of SAM. Although, more thorough experiments are needed to understand the utility of RGD on top of SAM.
Table 11: Comparison of RGD with SAM. Accuracy performance numbers have been reported over an average of 5 seeds.
Algorithm SGD SAM SAM + RGD
Accuracy 96.80 $\pm$ 0.04 97.31 $\pm$ 0.05 97.41 $\pm$ 0.09
D.4 Class Imbalance Experiments
@@
This section briefly discusses additional results from our experiments on the Class Imbalance domain with imbalanced CIFAR-10 and CIFAR-100 datasets. It is well known that DRO outputs models with good tail performance (Duchi & Namkoong, 2018). Since RGD directly solves the DRO objective, our models are also naturally endowed with this property. To demonstrate this, we extend our experiments on linear regression to a more realistic image dataset, where some classes appear very rarely in the data set while some appear very frequently. We use the Long-Tailed CIFAR dataset, where we reduce the number of training samples per class according to an exponential function as proposed by Cui et al. (2019). We define the imbalance factor of a dataset as the number of training samples in the largest class divided by the smallest. Similar to the works of Shu et al. (2019), we use a ResNet-32 architecture for training. Apart from Cross Entropy loss, we also include Focal Loss (Lin et al., 2017) and Class Balanced Loss (Cui et al., 2019) as additional baselines. We also experimented with the long-tailed CIFAR-100 dataset and showed that our proposed approach could again show significant improvements. Figure 5 illustrates the performance of our approach in comparison to other state-of-the-art methods. Overall, in comparison to the SOTA approach in this task (Class Balanced Loss), our proposed approach brings about an improvement of +0.79%. A more comprehensive comparison with additional state-of-the-art baselines such as L2RW (Ren et al., 2018), and Meta-Weight-Net (Shu et al., 2019) is illustrated in Table 13. Although these models use additional data as a meta-validation-set, our proposed approach outperforms L2RW and is roughly competitive with the Meta-Weight-Net model.
@@
Table 12 depicts the accuracy metric of models on various levels of the imbalance factor. From Table 12, we show that our proposed approach RGD outperforms other baselines such as Focal Loss and Class Balanced Loss by +0.79%. Furthermore, when models are trained on additional data, either by fine-tuning or by using a meta-learning framework to learn weights (such as Meta-Weight-Net and L2RW), we show that our proposed approach is competitively similar (-0.22%). Table 13 illustrates this analysis further. The performance metrics of the baseline approaches were taken from Shu et al. (2019). Additional comparisons against other losses, such as cross-entropy with label smoothing and large margin softmax loss, are shown in Table 14.
@@
Focal Loss Lin et al. (2017) 65.29 70.38 76.71 82.76 86.66 93.03 79.14 35.62 38.41 44.32 51.95 55.78 70.52 49.43
Class Balanced Loss Cui et al. (2019) 68.89 74.57 79.27 84.36 87.49 92.89 81.25 36.23 39.60 45.32 52.59 57.99 70.50 50.21
@@
L2RW Ren et al. (2018) $\star$ 66.51 74.16 78.93 82.12 85.19 89.25 77.69 33.38 40.23 44.44 51.64 53.73 64.11 47.92
Meta-Weight-Net Shu et al. (2019) $\star$ 68.91 75.21 80.06 84.94 87.84 92.66 81.60 37.91 42.09 46.74 54.37 58.46 70.37 51.65
@@
D.5 Vanilla Classification
This section briefly discusses a few additional results from our experiments on standard supervised learning (in particular vanilla classification). Table 15 depicts the performance of our other variant RGD-RevKL in comparison to the baseline approach.
EfficientNet finetuning.
We also show fine-tuning improvements of EfficientNet-v2-l over various tasks such as Cars and Food101 as depicted in Table 16. In these experiments, we take a pre-trained EfficientNet backbone and fine-tune it for various tasks.
DieT-S for ImageNet-1K classification.
Similarly, we also report that RGD boosts the performance of the baseline DeiT-S model by $0.1\%$ when trained from scratch on the Imagenet-1K benchmark as depicted in Table 16. Note that similar to our setup in ViT-S on Imagenet-1K benchmark, we perform no tuning, and simply use $\tau=1$, and same learning rate as baseline.
MLP for classification.
We also demonstrate that our proposed approach is simple and shows significant improvements, not only for SOTA approaches but also basic MLP procedures as depicted in Table 18, Table 19, and Table 20. These tables help showcase that the simple addition of our proposed approach does show significant improvements of +2.77% (in accuracy) on multi-class and +1.77% (in AUROC) on binary class tasks respectively.
bert-base-uncased MNLI QQP QNLI SST-2 MRPC RTE COLA Avg
Default 81.33 89.62 87.93 90.63 89.55 67.19 54.53 80.11
RGD-RevKL (Ours) 82.97 89.87 90.79 91.28 88.54 71.23 59.28 81.97
RGD (Ours) 83.06 91.06 90.35 91.78 88.28 71.48 58.56 82.05
Table 16: Ablation study to showcase the gain achieved by using RGD for various tasks in Vision benchmarks for ViT-S on Imagenet-1K, DeiT-S on Imagenet-1K and EfficientNet-v2-l on Finetuning tasks.
Imagenet-1K (ViT-S) Imagenet-1K (DeiT-S) Cars-FineTuning Food101-FineTuning
Default 78.1 80.05 92.03 92.65
RGD (Ours) 79.0 80.13 92.62 92.75
MiniGPT Pre-training.
We extend our work to large-scale tasks in NLP such as LLM pre-training which has become more prevalent over the recent years. Our experiments on miniGPT pre-training, as illustrated in this section, showcase the efficacy of our approach in these settings. miniGPT (Zhu et al., 2024) is a minimal implementation of a decoder-only transformer language model. We consider a 6 layer model and train on the lm1B small dataset which has 1B tokens. We trained for 100K steps with a batch size of 256. We used the default learning rate of 0.0016 for the baseline. For RGD, we fix the clipping threshold to 1 and tune the learning rate. We achieved 1% improvement on the eval log-perplexity score. Table 17 illustrates our results in this setting.
Table 17: Results on standard Lm1b pre-training task with miniGPT. Off-the-hat addition of RGD  leads to improvements of evaluation log-perplexity by $1\%$.
Dataset Lm1b eval Perplexity
Default 2.9218
RGD (Ours) 2.8979
D.6 Tabular Classification
This section discusses a few additional results from our experiments on Tabular classification. Table 21 depicts our proposed approach’s accuracy compared to other baselines on multi-class tabular datasets. Our method outperforms previous SOTA in this problem by +1.27%. Furthermore, Table 22 illustrates the AUROC score of our proposed approach in comparison to state-of-the-art baselines on binary-class tabular datasets. Our approach shows an improvement of +1.5% in this setting as well. The performance metrics of the baseline approaches were taken from Majmundar et al. (2022).
@@ D.6 Tabular Classification
RF Breiman (2001) 88.43 42.73 97.62 71.37 75.04
GBDT Friedman (2001) 88.71 45.7 100 72.96 76.84
RF-G Rahimi & Recht (2008) 89.84 29.32 97.65 71.57 72.10
MET-R Majmundar et al. (2022) 88.84 28.94 97.44 69.68 71.23
VIME Yoon et al. (2020) 80.36 34.00 95.77 62.80 68.23
DACL+ Verma et al. (2021) 81.40 39.70 91.40 64.23 69.18
SubTab Ucar et al. (2021) 87.59 39.34 98.31 42.36 66.90
TabNet Arik & Pfister (2021) 88.18 33.75 96.63 65.13 70.92
MET Majmundar et al. (2022) 91.68 47.82 99.19 76.71 78.85
@@
RF Breiman (2001) 64.36 91.53 77.57 99.62 83.27
GBDT Friedman (2001) 64.4 92.5 78.77 99.34 83.75
RF-G Rahimi & Recht (2008) 54.45 90.09 80.32 52.65 69.37
MET-R Majmundar et al. (2022) 53.2 83.54 79.17 82.03 74.49
VIME Yoon et al. (2020) 57.27 87.37 74.28 94.87 78.45
DACL+ Verma et al. (2021) 61.18 89.01 75.32 86.63 78.04
SubTab Ucar et al. (2021) 64.92 88.95 76.57 88.93 79.00
TabNet Arik & Pfister (2021) 69.40 77.30 80.91 96.98 81.15
@@
D.7 Meta-Learning
This section discusses some additional results from our experiments in the meta-learning domain. Table 23 depicts a complete table and comparison of our proposed approach on the MAML baseline compared to others. Overall, we notice improvement across the board, especially in the outliers, as shown in the Worst-kmetrics. Note that although our RGD has been applied on MAML (Finn et al., 2017) in our current experiments, our approach is analogous to the model and can be extended to other meta-learning techniques such as Protonet (Snell et al., 2017), CNAPs (Requeima et al., 2019), etc. as well. The performance metrics of the baseline approaches were taken from (Kumar et al., 2023).
@@ D.8 DomainBed
ERM Gulrajani & Lopez-Paz (2021) 85.5 $\pm$ 0.1 77.5 $\pm$ 0.4 66.5 $\pm$ 0.2 40.9 $\pm$ 0.1 67.6
IRM Arjovsky et al. (2019) 83.5 $\pm$ 0.8 78.5 $\pm$ 0.5 64.3 $\pm$ 2.2 33.9 $\pm$ 2.8 65.1
GroupDRO Sagawa* et al. (2020) 84.4 $\pm$ 0.8 76.7 $\pm$ 0.6 66.0 $\pm$ 0.7 33.3 $\pm$ 0.2 65.1
Mixup Yan et al. (2020a) 84.6 $\pm$ 0.6 77.4 $\pm$ 0.6 68.1 $\pm$ 0.3 39.2 $\pm$ 0.1 67.33
MLDG Li et al. (2018a) 84.9 $\pm$ 1.0 77.2 $\pm$ 0.4 66.8 $\pm$ 0.6 41.2 $\pm$ 0.1 67.53
CORAL Sun & Saenko (2016) 86.2$\pm$ 0.3 78.8$\pm$ 0.6 68.7$\pm$ 0.3 41.5 $\pm$ 0.1 68.8
MMD Li et al. (2018b) 84.6 $\pm$ 0.5 77.5 $\pm$ 0.9 66.3 $\pm$ 0.1 23.4 $\pm$ 9.5 62.95
DANN Ganin et al. (2016) 83.6 $\pm$ 0.4 78.6 $\pm$ 0.4 65.9 $\pm$ 0.6 38.3 $\pm$ 0.1 66.6
CDANN Li et al. (2018c) 82.6 $\pm$ 0.9 77.5 $\pm$ 0.1 65.8 $\pm$ 1.3 38.3 $\pm$ 0.3 66.05
MTL Blanchard et al. (2021) 84.6 $\pm$ 0.5 77.2 $\pm$ 0.4 66.4 $\pm$ 0.5 40.6 $\pm$ 0.1 67.2
SagNet Nam et al. (2021) 86.3 $\pm$ 0.2 77.8 $\pm$ 0.5 68.1 $\pm$ 0.1 40.3 $\pm$ 0.1 68.13
ARM Zhang et al. (2021) 85.1 $\pm$ 0.4 77.6 $\pm$ 0.3 64.8 $\pm$ 0.3 35.5 $\pm$ 0.2 65.75
VREx Krueger et al. (2021) 84.9 $\pm$ 0.6 78.3 $\pm$ 0.2 66.4 $\pm$ 0.6 33.6 $\pm$ 2.9 65.8
RSC Huang et al. (2020) 85.2 $\pm$ 0.9 77.1 $\pm$ 0.5 65.5 $\pm$ 0.9 38.9 $\pm$ 0.5 66.68
MIRO Cha et al. (2022) 85.4 $\pm$ 0.4 79.0 $\pm$ 0.0 70.5 $\pm$ 0.4 44.3 $\pm$ 0.2 69.8
@@
D.8.1 DomainBed Benchmark
D.8.2 Additional Results
In this section, we briefly discuss additional results from our DomainBed experiments. Table 24 depicts a complete table and comparison of our proposed approach to a multitude of state-of-the-art approaches in this field. Furthermore, we also show that our proposed approach outperforms previous SOTA by +0.7%.
Moreover, we also report the performance improvements when RGD is trained with model weight averaging methods such as SWAD Cha et al. (2021). Table 29 depicts the performance improvements of RGD over SWAD.
Table 29: Results on DomainBed (Model selection: training-domain validation set) on the model weight averaging models such as SWAD Cha et al. (2021): The bottom partition shows results of our method with RGD loss.
Algorithm PACS VLCS OfficeHome DomainNet Avg.
SWAD
Default Cha et al. (2021) 86.5 $\pm$ 1.0 76.0 $\pm$ 0.7 66.3 $\pm$ 0.2 43.8 $\pm$ 0.1 68.15
RGD (Ours) 87.6 $\pm$ 0.2 75.4 $\pm$ 1.1 67.5 $\pm$ 0.3 44.0 $\pm$ 0.1 68.63
Furthermore, we also present the per-environment breakdown of our approach in various datasets in Table 25, Table 26, Table 27, and Table 28 for PACS, VLCS, OfficeHome, and DomainNet respectively. The performance metrics of the baseline approaches were taken from Gulrajani & Lopez-Paz (2021).
D.9 Convergence of RGD and additional costs